OpenStudy (anonymous):

Find an integer n whose first digit is three, such that 3n/2 is obtained by removing the 3 at the beginning and putting it at the end.

OpenStudy (anonymous):

\[n=3a,3/2n=a3\]

OpenStudy (anonymous):

@TuringTest @dpaInc

OpenStudy (cruffo):

The number is 3529411764705882 I started by looking for the last digit of n. n = 3....k and 3n/2 = .....3 3*k/2 = 3 means k has to be 2. n = 3 ... 2 and 3n/2 = .... 23 Step 1 .....? x 3 ----- 3 ? = 1 Step 2 1 __________ 2 | 3 ........ 2 | v _____ 0 2 2 ---- 0 Step 3: ... ?1 x 3 ----- 23 ? = 4 carry 1 Step 4: 41 __________ 2 | 3 ........?2 2 goes into ? 4 times with no remainder | | v | 08 | 8 v ----- 0 2 2 ---- 0 ? = 8 so n = 3 ... 82 and 3n/2 = ...823 Step 5: 1 ... ?41 x 3 ------ 823 ? = 9 since 3 * 9 + 1 = 28 (carry 2) and continue... No joking, this took a while and I started to think I wasn't getting anywhere, but it did finally work out.

OpenStudy (anonymous):

medal medal medal